OpenStudy (anonymous):

Write an equation AND solve for the number: 16 increased by a number is -20. Find the number.

OpenStudy (aroub):

16+x=-20 x=-16-20 x=-36

OpenStudy (anonymous):

16 + x = -20 x=-36.00000000000000000000000000000000000

OpenStudy (anonymous):

theres one more 0 at the end of that ryan!!

OpenStudy (anonymous):

You have an equation as follows: x + 16 = -20 This is because we are unsure of the number being increased by 16. The solution is as follows: x = -20 - 16 x = -36 Now you can plug in and work it to check your answer: -36 + 16 = -20 Check! It worked. There ya go! Hope this helps. If it does, please choose good answer. Visit http://www.tutorsean.net for free math videos.